# Evaluating The Explainability of State-of-the-Art Deep Learning-based Network Intrusion Detection Systems

### v3

@@ V Experimental Results
NIDSs Evaluated: We have used all the state-of-the-art DL-based NIDSs presented in Section III-C (Kitsune, HorusEeye, ENIDrift and HELAD) for our analysis. When we refer to HorusEye from this point onwards, we are referring to its Magnifier component which performs deep learning-based anomaly detection.
@@
Datasets: We use the Kitsune Mirai dataset [5] and the CICIDS-2017 dataset [17] for training/testing the DL-based NIDSs under consideration. The Mirai PCAP trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. We use the Kitsune code to extract features from the PCAP trace through the DIS method. The CICIDS-2017 dataset [17] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. The CICIDS-2017 dataset provides PCAP traces, labeled network traffic flows and CSV files consisting of pre-selected features extracted from traffic flows. We use the raw PCAP traces directly and extract features from them using the Kitsune DIS method to maintain consistency. Below, we present the results using Mirai dataset only. Since the results obtained with CICIDS-2017 dataset are similar to those obtained with the Mirai dataset, we include them in the Appendix Section A.
@@
V-A Case Study- Kitsune $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table I. The sample size refers to the fraction of the training dataset to use to train the student decision tree model. Fidelity is defined as the R-squared value between Kitsune’s predictions and those obtained by the DT explanation. Using TRUSTEE with $30\%$ of the original Mirai dataset samples and no pruning results in the DT explanation which achieves 0.786 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 1 and achieves 0.697 fidelity compared to Kitsune.
$S=\alpha*(F1-score)+(1-\alpha)*(Fidelity),0\leq\alpha\leq 1$ (5)
  • $HpHp\_0.01\_pcc\_0\_1$, $HpHp\_0.1\_pcc\_0\_1$- Correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time windows 0.01 (1 minute) and 0.1 (10 seconds)
  • $HpHp\_0.1\_weight\_0$- Weights (packet count) aggregated by traffic sent between a set of source and destination IP addresses and associated with the time window 0.1 (1.5 seconds).
Thus, Kitsune’s DT relies mainly on the volume of packets and sizes of packets exchanged per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 133377, 70, 66689 0.786
30% Yes (k=3) 19, 9, 10 0.697
30% Yes (k=10) 37, 13, 19 0.721
TABLE I: TRUSTEE analysis results for Kitsune Mirai model
The base value is $0.0646$. Features such as $HpHp\_0.1\_weight\_0$ and $MI\_dir\_0.1\_weight$ have a negative impact on the prediction value. $HpHp\_0.1\_weight\_0$ is the most crucial feature, as the contribution has a broader range. The total negative contribution is greater than the positive contribution, and the final predicted value is lesser than the base value. As a result, the class is predicted as benign. The feature $HpHp\_0.1\_weight\_0$ corresponds to weights aggregated by traffic sent between a set of source and destination IP addresses using a specific protocol (ARP/TCP/UDP) with time window 0.1 (10 seconds) while the feature $MI\_dir\_0.1\_weight$ corresponds to weights aggregated by source MAC and IP addresses, with time window 0.1 (10 seconds).
Using kernel SHAP analysis, Fig. 3 shows a summary of how the top features in Mirai dataset impact the Kitsune NIDS model’s output through a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 1) which takes $625.70$ seconds on an average. Running SHAP analysis for the whole dataset will take substantial amount of time and computing resources. Each instance of the given explanation is represented by a single dot on each feature row. The $x$ position of the dot is determined by the SHAP value of that feature, and dots pile up along each feature row to show density. Color is used to display the original value of a feature. The features are ordered using the mean absolute value of the SHAP values for each feature. Features such as $MI\_dir\_0.1\_weight$, $HH\_0.01\_weight\_0$ and $HH\_0.1\_weight\_0$ have the most impact on the prediction value, in that order. $HH\_0.01\_weight\_0$ and $HH\_0.1\_weight\_0$ correspond to weights aggregated by traffic sent between a set of source and destination IP addresses with time windows 0.01 (1 minute) and 0.1 (10 seconds) respectively.
Figure 3: SHAP beeswarm plot for Kitsune Mirai model
Detection of Inductive Bias: It was shown in [7] that Kitsune NIDS is vulnerable to variable packet rate samples when used with the Mirai dataset introduced in [5]. The authors evaluated the following threshold ranges (in packets per second) for ARP requests originating from devices infected with Mirai: (1) 10-50, (2) 30-70 and (3) 50-90. The original trace as well as the tampered traces are shown in Fig. 4. It can be observed that tampering had a significant effect on the difference between benign and attack traffic in terms of number of packets per second. Kitsune was executed on each trace with an identical number of training samples. The results showed that the RMSE values generated by Kitsune NIDS decreased significantly with a drop in the attack traffic volume, reaching the values corresponding to benign traffic.
Figure 4: Number of packets per second for original Kitsune Mirai trace and tampered traces. Blue portions of the plot correspond to benign traffic while red portions correspond to malicious traffic inclusive of benign and attack traffic.
V-B Case Study- HorusEye $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table II. Using TRUSTEE with $30\%$ of the Mirai dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-1.365$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-1.038e-06$) compared to original HorusEye model. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
  1. 1.
    $HH\_3\_pcc\_0\_1$- Correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 3 (500 milli-seconds).
  2. 2.
    $HH\_1\_std\_0$- Standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time window 1 (1.5 seconds).
  3. 3.
    $MI\_dir\_0.01\_mean$- Mean of packet sizes aggregated by source MAC and IP addresses, with time window 0.01 (1 minute).
Thus, HorusEye’s DT relies mainly on the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 45, 11, 23 -1.365
30% Yes (k=10) 1, 0, 1 -1.038e-06
TABLE II: TRUSTEE analysis results for HorusEye Mirai model
SHAP Analysis: Using kernel SHAP analysis, Fig. 6 shows a summary of how the top features in Mirai dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 5). Features such as $HH\_1\_std\_0$, $HH\_5\_std\_0$ and $HpHp\_0.01\_magnitude\_0\_1$ have the most impact on the prediction value, in that order. $HH\_1\_std\_0$ and $HH\_5\_std\_0$ correspond to the standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time windows 1 (1.5 seconds) and 5 (100 ms) respectively, while $HpHp\_0.01\_magnitude\_0\_1$ corresponds to the absolute magnitude of two packet size streams aggregated by the traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
Figure 6: SHAP beeswarm plot for HorusEye Mirai model
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HorusEye was executed on each trace. In Fig. 7(a), we show the results for first 200,000 packets of each traces. In Fig. 7(b), we also compare the expected RMSE (produced by HorusEye in the original trace) and the predicted RMSE for each tampered trace. The results show that the RMSE values generated by HorusEye decrease with a drop in the attack traffic volume. However, the difference between the RMSE values produced by the original and tampered traces is not as pronounced as in the case of Kitsune NIDS. This experiment demonstrates that the Mirai use case from HorusEye is vulnerable to variable packet rate samples, though to a lesser extent than Kitsune. It should be noted here that HorusEye [28] authors had recognized this issue (without an empirical analysis though) by stating that the frequency of Mirai attacks is too high in the original dataset to reflect real botnet infection. Therefore, the authors had randomly down-sampled the Mirai attack traffic in Kitsune dataset to consist of 40,000 packets during HorusEye’s performance evaluation.
(a) Predicted RMSE vs packet count
(b) Predicted vs expected RMSE (produced by HorusEye in the original trace)
Figure 7: HorusEye RMSE values for the first 200k packets from original and tampered traces.
V-C Case Study- ENIDrift $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table III. Using TRUSTEE with $30\%$ of the Mirai dataset samples and no pruning results in the DT explanation which achieves 0.715 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 8) which achieves negative fidelity ($\sim-1.536e-09$) compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are $f44$ and $f9$. ENIDrift does not extract explicit features from incoming packets. Instead, it uses ip2V incremental embedding technique based to embed packets to vectors. The embedding vector length in ENIDrift implementation is 200.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 5, 2, 3 0.715
30% Yes (k=10) 1, 0, 1 -1.536e-09
TABLE III: TRUSTEE analysis results for ENIDrift Mirai model
Figure 9: SHAP beeswarm plot for ENIDrift Mirai model
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and ENIDrift was executed on each trace. In Fig. 10(a), we show the results for first 120,000 packets of each traces. In Fig. 10(b), we also compare the expected (weighted) ensemble scores (produced by ENIDrift in the original trace) and the predicted (weighted) ensemble scores for each tampered trace. The results show that the ensemble scores produced by ENIDrift do not have a discernable pattern of dependence on the volume of the attack traffic encountered. Since the distance of the dots for tampered traces from the diagonal line in Fig. 12(b) does not follow any pattern, the effect that tampering with the Mirai attack traces had on the predicted ensemble scores and hence, whether the Mirai use case from ENIDrift is vulnerable to variable packet rate samples or not, remains inconclusive.
(a) Predicted ensemble score vs packet count
(b) Predicted vs expected ensemble score (produced by ENIDrift in the original trace)
Figure 10: ENIDrift ensemble scores for the first 120k packets from original and tampered traces.
V-D Case Study- HELAD $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table IV. Using TRUSTEE with $30\%$ of the Mirai dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 11) which achieves 0.0 fidelity compared to HELAD. A fidelity value of 0.0 means that the DT explanation fits HELAD predictions only as good as the mean of the predictions.
  1. 1.
    $MI\_dir\_3.0\_weight$- Weights aggregated by source MAC and IP addresses, with time window 3.0 (500 milli-seconds).
  2. 2.
    $HH\_1.0\_magnitude\_0\_1$- Absolute magnitude of two packet size streams aggregated by the traffic sent between a set of source and destination IP addresses with time window 1.0 (1.5 seconds).
  3. 3.
    $HH\_0.01\_weight\_0$- Weights aggregated by traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
Thus, HELAD’s DT relies mainly on the volume of packets and the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size Top-k pruning used? DT size, depth, leaves Fidelity
30% No 26289, 100, 13145 0.0
30% Yes (k=10) 435, 94, 218 0.0
TABLE IV: TRUSTEE analysis results for HELAD Mirai model
Detection of Inductive Bias: Following a similar approach as the Kitsune case, three separate tampered traces with different attack traffic volumes were obtained from the Mirai dataset and HELAD was executed on each trace. In Fig. 12(a), we show the results for first 200,000 packets of each traces. In Fig. 12(b), we also compare the expected anomaly scores (produced by HELAD in the original trace) and the predicted anomaly scores for each tampered trace. The results show that the anomaly scores produced by HELAD for the original and tampered traces overlap with each other and therefore, the scores do not have a significant dependence on the volume of the attack traffic encountered. Since the dots for tampered traces are close to the diagonal line in Fig. 12(b), tampering with the Mirai attack traces had limited outcome on the predicted anomaly scores. This experiment demonstrates that the Mirai use case from HELAD is not vulnerable to variable packet rate samples to a significant extent.
(a) Predicted anomaly score vs packet count
(b) Predicted vs expected anomaly score (produced by HELAD in the original trace)
Figure 12: HELAD anomaly scores for the first 200k packets from original and tampered traces.
V-E Comparative Analysis
In this sub-section, we use xAI evaluation metrics to quantify the effectiveness of the explanations generated for the DL-based NIDSs under consideration by TRUSTEE and kernel SHAP. We use those metrics to compare the explanations across NIDSs. Currently, the existing implementations of objective xAI evaluation metrics as explained in Section III-B support kernel SHAP only and none of the implementations support TRUSTEE. TRUSTEE’s implementation comes with its own metric called fidelity which quantifies how accurate the TRUSTEE DT is with respect to the original black-box ML model. We also use the TRUSTEE-SHAP agreement scores proposed in Section IV-B to compare the level of agreement between TRUSTEE DT and kernel SHAP explanations generated for the DL-based NIDSs considered.
@@ V-E Comparative Analysis
Using the CICIDS-2017 dataset, if we compare the fidelity values obtained for the NIDSs considered (refer Appendix Section A), without any pruning, the highest fidelity values are obtained for ENIDrift (0.929) out of the four NIDSs. With top-k pruning ($k=10$), the highest fidelity values are obtained for Kitsune (0.919) out of the four NIDSs. Since HorusEye and HELAD NIDSs yield negative and zero fidelity values, TRUSTEE DT explanations do not serve much purpose. With top-k pruning, TRUSTEE is most useful for explaining the decisions of Kitsune NIDS followed by ENIDrift NIDS. Across the datasets, with top-k pruning, Kitsune NIDS yields the highest fidelity values. The comparison of fidelity values obtained during TRUSTEE analysis of the DL-based NIDSs considered in this paper is shown in Table V.
NIDS Name Fidelity (Without pruning) Fidelity (with top-k pruning, k=10)
Mirai dataset CICIDS-2017 dataset Mirai dataset CICIDS-2017 dataset
@@
TABLE V: Comparison of TRUSTEE fidelity values with and without pruning for DL-based NIDSs considered
@@
Level of TRUSTEE-SHAP Agreement: We set the value of $m=3$ for the $m$-point average TRUSTEE-SHAP agreement score. Each of the data subsets used for calculation of TRUSTEE-SHAP agreement scores were required to have a minimum number of 300 samples. The results are shown in Table VI. HorusEye, with a score of $0.381$, exhibits the highest level of TRUSTEE-SHAP agreement, followed by Kitsune, with a score of $0.167$ while ENIDrift has a score of $0.0$. We also find that the TRUSTEE-SHAP agreement scores obtained for an NIDS for the different data subsets vary significantly. For example, for Kitsune, the scores obtained for three different data subsets were $0.4$, $0.6$ and $0.1$. This supports our earlier assumption that the level of agreement between TRUSTEE DT explanation and SHAP explanation for a blackbox NIDS model may depend on the specific data subset which is used to conduct the TRUSTEE-SHAP comparison.
NIDS $m$-point avg. TRUSTEE-SHAP agreement score (m=3)
@@
TABLE VI: Comparison of $m$-point average TRUSTEE-SHAP agreement scores for DL-based SOTA NIDSs trained with Kitsune Mirai dataset
VI Discussion and Future Work

### v4

@@ 4 Experimental Results
NIDSs Evaluated: We have used the state-of-the-art DL-based NIDSs presented in Section 2.3 (Kitsune, HorusEeye, ENIDrift and HELAD) for our analysis. When we refer to HorusEye from this point onwards, we are referring to its Magnifier component which performs deep learning-based anomaly detection. Since the feature extraction algorithm code for Deep ResNIDS implementation is not publicly available, we have not included it in our analysis.
@@
Datasets: We use the Kitsune Mirai dataset [5] and the CICIDS-2017 dataset [28] for training/testing the DL-based NIDSs under consideration. The Mirai PCAP trace consists of $\sim$ 760k packets of synthetically generated attack in a network with nine IoT devices, first $\sim$120k packets consist of benign traffic, remaining $\sim$640k packets have anomalous traffic. We use the Kitsune code to extract features from the PCAP trace through the DIS method. The CICIDS-2017 dataset [28] consists of five days of packet capture (July 3 to July 7, 2017) from a small-scale enterprise-like network, with the first day consisting of benign background traffic and the other days consisting of attack traffic (brute force SSH, DoS, Heartbleed, web attack, infiltration, etc.). It consists of significant drift over the data collection period in terms of different attack tools used. The CICIDS-2017 dataset provides PCAP traces, labeled network traffic flows and CSV files consisting of pre-selected features extracted from traffic flows. We use the raw PCAP traces directly and extract features from them using the Kitsune DIS method to maintain consistency. Below, we present the explanations generated using Mirai dataset only due to space limitations though we have included a few results obtained using CICIDS-2017 dataset in Section 4.1.
@@
4.1 Comparative Analysis
In this sub-section, we use XAI evaluation metrics to quantify the effectiveness of the explanations generated for the DL-based NIDSs under consideration by TRUSTEE and kernel SHAP. TRUSTEE’s implementation comes with its own metric called fidelity which quantifies how accurate the TRUSTEE DT is with respect to the original black-box ML model. We also use the TRUSTEE-SHAP agreement score proposed in Section 3.2 to compare the level of agreement between TRUSTEE DT and kernel SHAP explanations generated for the DL-based NIDSs considered and security-focused criteria such as adversarial robustness/stability to compare TRUSTEE versus SHAP explanations.
@@ 4.1 Comparative Analysis
Using the CICIDS-2017 dataset, if we compare the fidelity values obtained for the NIDSs considered without any pruning, the highest fidelity values are obtained for ENIDrift (0.929) out of the four NIDSs. With top-k pruning ($k=10$), the highest fidelity values are obtained for Kitsune (0.919) out of the four NIDSs. Since HorusEye and HELAD NIDSs yield negative and zero fidelity values, TRUSTEE DT explanations do not serve much purpose. With top-k pruning, TRUSTEE is most useful for explaining the decisions of Kitsune NIDS followed by ENIDrift NIDS. The comparison of fidelity values obtained during TRUSTEE analysis of the DL-based NIDSs considered in this paper is shown in Table 1.
Takeaway 1: Across the datasets, with top-k pruning, Kitsune NIDS yields the highest TRUSTEE DT fidelity values.
NIDS Name
Fidelity
(Without pruning)
Fidelity (with top-k
pruning, k=10)
Mirai
dataset
CICIDS-2017
dataset
Mirai
dataset
CICIDS-2017
dataset
@@
Table 1: Comparison of TRUSTEE fidelity values with and without pruning for DL-based NIDSs considered
@@
Level of TRUSTEE-SHAP Agreement: We set the value of $m=3$ for the $m$-point average TRUSTEE-SHAP agreement score. Each of the data subsets used for calculation of TRUSTEE-SHAP agreement scores were required to have a minimum number of 300 samples. The results are shown in Table 2. HorusEye, with a score of $0.381$, exhibits the highest level of TRUSTEE-SHAP agreement, followed by Kitsune, with a score of $0.167$ while ENIDrift has a score of $0.0$. We also find that the TRUSTEE-SHAP agreement scores obtained for an NIDS for the different data subsets vary significantly. For example, for Kitsune, the scores obtained for three different data subsets were $0.4$, $0.6$ and $0.1$.
Takeaway 2: The level of agreement between TRUSTEE DT explanation and SHAP explanation for a blackbox NIDS model may depend on the specific data subset which is used to conduct the TRUSTEE-SHAP comparison.
NIDS
$m$-point avg. TRUSTEE-SHAP
agreement score (m=3)
@@
Table 2: Comparison of $m$-point average TRUSTEE-SHAP agreement scores for DL-based SOTA NIDSs trained with Kitsune Mirai dataset
Adversarial Robustness: SHAP explanations have already been shown to be vulnerable to adversarial input perturbations as per [25]. Further, global explanation methods such as Partial Dependence (PD) plots have also been subject to adversarial attacks either by poisoning the dataset used to obtain the PD explanation for a biased black-box model to manipulate the explanation [29] or using extrapolated permuted PD data [30] to preserve the original black-box model predictions while hiding the model’s discriminatory behaviour. There is no existing work on adversarial attacks against TRUSTEE (also a global explanation method), though we surmise possible adversarial attacks which poison the dataset (used to obtain a biased black-box model’s explanation) in a way that the curated dataset (sampled from the poisoned dataset) used to guide the surrogate DT model, can be used to manipulate the final DT explanation while maintaining high fidelity.
Takeaway 3: SHAP and TRUSTEE explanations are both vulnerable to adversarial attacks, though further work is needed to quantify their degree of robustness.
Stability: We evaluate the stability scores of TRUSTEE and SHAP explanations for each DL-based NIDS trained over Mirai dataset by calculating the average intersection size (as defined in Equation 6) of the top $10$ features in the explanations generated by each method over $3$ runs. The scores obtained are shown in Table 3. The average intersection size for ENIDrift is lowest among all the NIDSs for TRUSTEE and SHAP explanations. Across TRUSTEE and SHAP, HorusEye has the highest stability, closely followed by Kitsune. Overall, SHAP fares slightly better than TRUSTEE in terms of stability. Though kernel SHAP explanations for individual data instances often suffer from instability [31], we have used the mean absolute value of the SHAP values for each feature across all the data samples which improves the stability of resulting explanation.
Takeaway 4: Across the NIDSs, SHAP explanations yield better stability than the TRUSTEE explanations.
Method Kitsune HorusEye ENIDrift HELAD
SHAP 0.933 1.000 0.133 n/a
TRUSTEE 0.833 1.000 0.033 0.200
Table 3: Average intersection size between top $k=10$ features of explanations for 3 runs. Values closer to 1 indicate greater stability.
4.2 Case Study- Kitsune $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for Kitsune are shown in Table 4. The sample size refers to the fraction of the training dataset to use to train the student decision tree model. Fidelity is defined as the R-squared value between Kitsune’s predictions and those obtained by the DT explanation. Using TRUSTEE with $30\%$ of the original Mirai dataset samples and no pruning results in the DT explanation which achieves 0.786 fidelity compared to Kitsune. Using TRUSTEE with $30\%$ of the dataset samples and top-k pruning method (setting $k=3$) results in a DT explanation that is shown in Fig. 1 and achieves 0.697 fidelity compared to Kitsune. We now explain the decision tree briefly: when the feature HpHp_0.01_pcc_0_1HpHp\_0.01\_pcc\_0\_1 is less than or equal to the value $147439.125$, the control passes to the left branch, whereas when the opposite is true, the control passes to the right branch. Squared error refers to the mean-squared error between the target and predicted value for the data points covered in the tree node, samples refers to the number of data points covered in the node and value refers to the value of the target variable. Dark colored nodes with no feature names represent the leaf nodes.
$S=\alpha*(F1\ score)+(1-\alpha)*(Fidelity),0\leq\alpha\leq 1$ (7)
  • HpHp_0.01_pcc_0_1HpHp\_0.01\_pcc\_0\_1, HpHp_0.1_pcc_0_1HpHp\_0.1\_pcc\_0\_1- Correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time windows 0.01 (1 minute) and 0.1 (10 seconds)
  • HpHp_0.1_weight_0HpHp\_0.1\_weight\_0- Weights (packet count) aggregated by traffic sent between a set of source and destination IP addresses and associated with the time window 0.1 (1.5 seconds).
Takeaway 5: Kitsune’s DT relies mainly on the volume of packets and sizes of packets exchanged per time frame to determine if an attack is underway.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
30% No 133377, 70, 66689 0.786
30% Yes (k=3) 19, 9, 10 0.697
30% Yes (k=10) 37, 13, 19 0.721
Table 4: TRUSTEE analysis results for Kitsune Mirai model
The base value is $0.0646$. Features such as HpHp_0.1_weight_0HpHp\_0.1\_weight\_0 and MI_dir_0.1_weightMI\_dir\_0.1\_weight have a negative impact on the prediction value. HpHp_0.1_weight_0HpHp\_0.1\_weight\_0 is the most crucial feature, as the contribution has a broader range. The total negative contribution is greater than the positive contribution, and the final predicted value is lesser than the base value. As a result, the class is predicted as benign. The feature HpHp_0.1_weight_0HpHp\_0.1\_weight\_0 corresponds to weights aggregated by traffic sent between a set of source and destination IP addresses using a specific protocol (ARP/TCP/UDP) with time window 0.1 (10 seconds) while the feature MI_dir_0.1_weightMI\_dir\_0.1\_weight corresponds to weights aggregated by source MAC and IP addresses, with time window 0.1 (10 seconds).
Using kernel SHAP analysis, Fig. 3 shows a summary of how the top features in Mirai dataset impact the Kitsune NIDS model’s output through a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 1) which takes $625.70$ seconds on an average. Running SHAP analysis for the whole dataset will take substantial amount of time and computing resources. Each instance of the given explanation is represented by a single dot on each feature row. The $x$ position of the dot is determined by the SHAP value of that feature, and dots pile up along each feature row to show density. Color is used to display the original value of a feature. The features are ordered using the mean absolute value of the SHAP values for each feature. Features such as MI_dir_0.1_weightMI\_dir\_0.1\_weight, HH_0.01_weight_0HH\_0.01\_weight\_0 and HH_0.1_weight_0HH\_0.1\_weight\_0 have the most impact on the prediction value, in that order. HH_0.01_weight_0HH\_0.01\_weight\_0 and HH_0.1_weight_0HH\_0.1\_weight\_0 correspond to weights aggregated by traffic sent between a set of source and destination IP addresses with time windows 0.01 (1 minute) and 0.1 (10 seconds) respectively.
Figure 3: SHAP beeswarm plot for Kitsune Mirai model
4.3 Case Study- HorusEye $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for HorusEye are shown in Table 5. Using TRUSTEE with $30\%$ of the Mirai dataset samples with no pruning results in a DT explanation which achieves negative fidelity ($\sim-1.365$) compared to original HorusEye model. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation which achieves negative fidelity ($\sim-1.038e-06$) compared to original HorusEye model. Negative fidelity values mean that the DT explanation fits HorusEye predictions worse than the mean of the predictions.
  1. 1.
    HH_3_pcc_0_1HH\_3\_pcc\_0\_1- Correlation coefficient between two packet size streams aggregated by traffic sent between a set of source and destination IP addresses with time window 3 (500 milli-seconds).
  2. 2.
    HH_1_std_0HH\_1\_std\_0- Standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time window 1 (1.5 seconds).
  3. 3.
    MI_dir_0.01_meanMI\_dir\_0.01\_mean- Mean of packet sizes aggregated by source MAC and IP addresses, with time window 0.01 (1 minute).
Takeaway 6: HorusEye’s DT relies mainly on the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
30% No 45, 11, 23 -1.365
30% Yes (k=10) 1, 0, 1 -1.038e-06
Table 5: TRUSTEE analysis results for HorusEye Mirai model
SHAP Analysis: Using kernel SHAP analysis, Fig. 5 shows a summary of how the top features in Mirai dataset impact the HorusEye NIDS model’s output using a beeswarm plot for randomly selected 1000 data samples corresponding to the benign root node-leaf node path in the TRUSTEE DT (Fig. 4). Features such as HH_1_std_0HH\_1\_std\_0, HH_5_std_0HH\_5\_std\_0 and HpHp_0.01_magnitude_0_1HpHp\_0.01\_magnitude\_0\_1 have the most impact on the prediction value, in that order. HH_1_std_0HH\_1\_std\_0 and HH_5_std_0HH\_5\_std\_0 correspond to the standard deviation of packet sizes aggregated by the traffic sent between a set of source and destination IP addresses with time windows 1 (1.5 seconds) and 5 (100 ms) respectively, while HpHp_0.01_magnitude_0_1HpHp\_0.01\_magnitude\_0\_1 corresponds to the absolute magnitude of two packet size streams aggregated by the traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
Figure 5: SHAP beeswarm plot for HorusEye Mirai model
4.4 Case Study- ENIDrift $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for ENIDrift are shown in Table 6. Using TRUSTEE with $30\%$ of the Mirai dataset samples and no pruning results in the DT explanation which achieves 0.715 fidelity compared to ENIDrift. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 6) which achieves negative fidelity ($\sim-1.536e-09$) compared to ENIDrift. This is expected since pruning the branches of a DT reduces its fidelity. The most prominent features ENIDrift uses to determine anomaly are $f44$ and $f9$. ENIDrift does not extract explicit features from incoming packets. Instead, it uses ip2V incremental embedding technique based to embed packets to vectors. The embedding vector length in ENIDrift implementation is 200.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
30% No 5, 2, 3 0.715
30% Yes (k=10) 1, 0, 1 -1.536e-09
Table 6: TRUSTEE analysis results for ENIDrift Mirai model
Figure 7: SHAP beeswarm plot for ENIDrift Mirai model
4.5 Case Study- HELAD $\times$ Mirai
TRUSTEE Analysis: The results of TRUSTEE analysis for HELAD are shown in Table 7. Using TRUSTEE with $30\%$ of the Mirai dataset samples and no pruning results in the DT explanation which achieves 0.0 fidelity compared to HELAD. Using TRUSTEE with $30\%$ of the dataset samples and top-k Pruning method (setting $k=10$) results in a DT explanation (shown in Fig. 8) which achieves 0.0 fidelity compared to HELAD. A fidelity value of 0.0 means that the DT explanation fits HELAD predictions only as good as the mean of the predictions.
  1. 1.
    MI_dir_3.0_weightMI\_dir\_3.0\_weight- Weights aggregated by source MAC and IP addresses, with time window 3.0 (500 milli-seconds).
  2. 2.
    HH_1.0_magnitude_0_1HH\_1.0\_magnitude\_0\_1- Absolute magnitude of two packet size streams aggregated by the traffic sent between a set of source and destination IP addresses with time window 1.0 (1.5 seconds).
  3. 3.
    HH_0.01_weight_0HH\_0.01\_weight\_0- Weights aggregated by traffic sent between a set of source and destination IP addresses with time window 0.01 (1 minute).
Takeaway 7: HELAD’s DT relies mainly on the volume of packets and the sizes of packets exchanged between a pair of IP addresses per time frame to determine if an attack is underway.
Sample size
Top-k pruning
used?
DT size, depth, leaves Fidelity
30% No 26289, 100, 13145 0.0
30% Yes (k=10) 435, 94, 218 0.0
Table 7: TRUSTEE analysis results for HELAD Mirai model
5 Discussion and Future Work